OpenStudy (anonymous):

Solve the equation x − 6 = 11 A 5 B 17 C −17 D −5

OpenStudy (anonymous):

x-6=11 +6=+6

OpenStudy (anonymous):

what do you get

OpenStudy (anonymous):

c

OpenStudy (anonymous):

No what is 11+6

OpenStudy (anonymous):

sorry positive 17 my mistake

OpenStudy (anonymous):

B

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks!